OpenStudy (ccswims):

**WILL MEDAL** A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t^2 + 34t + 80 The average rate of change of f(t) from t = 5 seconds to t = 7 seconds is _____ feet per second.

OpenStudy (ccswims):

I just need my answer checked... I got -308

OpenStudy (mathmate):

Only the sign is correct. Did you use average rate of change = \(\dfrac{\Delta f(t)}{\Delta t}\) where \(\Delta t\) = 7-5=2 seconds?

OpenStudy (ccswims):

I did -466 - 150/7 - 5

OpenStudy (ccswims):

which got me -616/2, then I got -308

OpenStudy (mathmate):

ok, remember that when you deal with a fraction, all numerators and denominators have implied brackets, so you understand the concept, but have a little slip in calculating. It actually is \(\dfrac{\color{red}{(}-466-(-150)\color{red}{)}}{2}=(-466-(-150))/2=(-466+150)/2=?\)

OpenStudy (ccswims):

OH! Okay, so it's actually -158 ft/per sec

OpenStudy (mathmate):

Exactly! Well done!